OpenStudy (anonymous):

help! algebra 2! 1. 2+4+8+... to 20 terms a. -1,048,575 b. -2,097,150 c. 2,097,150 d. 3,145,725 2. Find a_1 for the geometric series in which S_n=4,882,812, r=-5, and n=10 a. -2 b. -3 c. 2 d. 3 *I got C for both questions, i am not sure if i am right though

OpenStudy (anonymous):

\[ 2+ 4+ 8 + \cdots +2^{20}\\ \]

OpenStudy (anonymous):

\[ \sum_{i=1}^n a^i=\frac{a \left(1-a^{20}\right)}{1-a} \]

OpenStudy (anonymous):

Take a=2, and n=20 compute the answer

OpenStudy (anonymous):

\[ \frac{2 \left(1-2^{20}\right)}{1-2}=2,097,150 \]

OpenStudy (anonymous):

ok what more do you need help with?

OpenStudy (anonymous):

I was wondering if the 2nd answer was right.